use the elimination method to solve the system of equations
a+8b=1
a+3b=-4

Answer by checkley79(3341)

a+8b=1 multiply by -1 & add.
a+3b=-4
-a-8b=-1 add.
----------------
-5b=-5
b=-5/-5
b=1 ans.
a+8b=1
a+1*8=1
a+8=1
a=1-8
a=-7 ans.
Proof:
-7+3*1=-4
-7+3=-4
-4=-4